Five thick dictionaries each 10cm thick and each weighting 25N,are resting side by side flat on a table 1m high.how much work would be done against gravity to stack one on top of the other.


1
Expert's answer
2021-06-09T08:06:08-0400

"h_t=1m"

"h_d=10cm =0.1m"

"n=5"

"P=25N"

"\text{Potential energy of dictionaries lying on the table.}"

"\text{It is assumed that the dictionary is homogeneous and the height of the center}"

"\text{of gravity is half the width of the dictionary:}"

"E_{p0}= n*P*(h_t+\frac{h_d}{2})"

"E_{p0}= 5*25*(1+0.05)=131.25J"

"\text{The center of gravity of the stack of dictionaries on the table:}"

"h_c=h_t+\frac{n*h_d}{2}"

"h_c=1+\frac{5*0.1}{2}=1.25m"

"E_{p1}=n*P*h_c"

"E_{p1}=5*25*1.25=156.25J"

"W= E_{p1}-E_{p0}"

"W= 156.25-131.25=25J"

"\text{Answer: }W=25J"